# On the Effectiveness of Proposed Techniques to Reduce Energy Consumption in RAG Systems: A Controlled Experiment

### v1

@@ 1. Introduction
Large language models (LLMs) as the most recent wave of machine learning (ML) have shown impressive performance in natural language processing tasks and beyond (chang_survey_2024). However, they also face several challenges, including hallucination (Ji_2023_Hallucination_nlp; xu2025hallucinationinevitableinnatelimitation), outdated knowledge, and opaque or untraceable reasoning processes (gao2024retrievalaugmentedgenerationlargelanguage). Retrieval-augmented generation (RAG) has emerged as a popular technique to address shortcomings by integrating knowledge from external databases into the generation process (Ji_2023_Hallucination_nlp).
@@
As ML-enabled systems are increasingly adopted, more attention is also spent on their carbon footprint (Luccioni_2024; Bashir2024Climate). Due to factors such as increased technology adoption, cryptocurrency trends, and the rising demand for artificial intelligence, global data center electricity consumption is projected to range between 620 and 1,050 TWh by 2026 (IEA2024Electricity). The growing popularity of RAG systems is adding to this consumption, as they are increasingly applied in diverse domains such as code generation (parvez2021retrievalaugmentedcodegeneration; zhou2022docprompting), question answering (hu-etal-2022-logical; huang-etal-2021-unseen-entity), and AI for science (wang2023retrievalbasedcontrollablemoleculegeneration). Their energy consumption and carbon footprint should therefore be a subject of concern.
@@
While research on the environmental impact of ML-enabled systems has received more interest in recent years (Verdecchia2023GreenAI; Gezici2022), the quality attribute (QA) of environmental sustainability has received little direct attention in RAG systems. Existing research predominantly emphasizes other QAs, such as accuracy and latency, while largely neglecting energy efficiency. For instance, wang2024searchingbestpracticesretrievalaugmented investigated various RAG configurations but focused primarily on performance metrics like accuracy and latency. Similarly, a_synthesis_green_tactics_ml_sys have proposed 30 green architectural tactics for generic ML-enabled systems, but the practical applicability of these tactics within the context of RAG systems remains largely unexplored. This underscores a clear research gap: studying how to reduce energy consumption as a core quality concern in RAG systems and systematically exploring trade-offs with other system attributes. While some techniques have been proposed to improve the resource utilization or latency of RAG systems, no systematic study has examined the energy efficiency impact of RAG techniques to provide guidance to practitioners about their usage.
@@
To close this gap, the primary objective of this research is to analyze the energy consumption of various techniques in RAG systems while also examining potential trade-offs with response latency and the accuracy of generated answers. We conducted a controlled experiment (wohlin12) involving five techniques, some tested under different parameter settings, resulting in a total of nine configurations that we compared to a baseline industrial RAG system. To the best of our knowledge, this is the first study to conduct an in-depth investigation into different energy techniques for RAG systems. Our findings provide valuable insights for future research on the environmental sustainability of such systems and offer guidance on navigating trade-offs between energy usage, latency, and answer accuracy.
@@ 2. Related Work
a_synthesis_green_tactics_ml_sys presented a synthesis of green architectural tactics for ML-enabled systems. They compiled a catalog of 30 tactics, derived from an extensive review of Green AI literature and validated through an expert focus group. These tactics span six categories, namely data-centric strategies, algorithm design, model optimization, model training, deployment, and management, aimed at reducing energy consumption and enhancing computational efficiency throughout the ML system lifecycle. While their work provides valuable, actionable guidance for building sustainable ML-enabled systems, the proposed tactics are not specifically tailored to RAG systems, and evidence for their effectiveness in this context is missing. In this study, we focus specifically on production-level RAG systems and empirically evaluate the impact of selected techniques on energy efficiency.
@@
Using experiments to demonstrate that current LLM agents and RAG systems consume substantial amounts of energy, wu2025addressingsustainableaitrilemma introduced the “Sustainable AI Trilemma”, highlighting the tensions between AI capability, negative environmental impact, and digital inequality. They reported that common optimization steps, e.g., query optimization or compression, drastically increase energy use with diminishing returns and that LLM-dependent methods consume orders of magnitude more energy than non-LLM alternatives. While their paper crucially highlights these issues and inefficiencies within RAG systems, it stops at diagnosis and does not propose or study specific green techniques as a solution. In contrast, our research focuses explicitly on evaluating proposed technique variations for designing green RAG systems.
@@
Current techniques targeting RAG systems primarily focus on improving other key performance metrics, such as response time or reducing computational resource usage. In several cases, this can also lead to reductions in energy consumption. For example, arefeen2024irag examined the effect of threshold k selection in chunk filtering within the iRAG framework, demonstrating that a balanced selection of k enhances query efficiency and reduces computational waste. While increasing k results in longer query processing and greater computational overhead due to processing more candidate chunks, effective chunk filtering mitigates this by minimizing unnecessary computation while preserving high recall. However, there remains a lack of broad empirical evidence on the effectiveness and trade-offs of various energy-saving techniques specifically designed for RAG systems, especially for realistic systems from an industry context.
@@
Similarly, Sakar_Emekci_2025 conducted a comprehensive evaluation of various RAG systems, identifying optimal configurations that balance critical performance metrics like response accuracy, token efficiency, runtime, and hardware utilization across diverse domains. Their findings indicate that Reciprocal RAG, which generates and ranks multiple query variations to resolve ambiguity, achieved the highest similarity score but at the cost of significantly higher token usage and longer run times. In contrast, the Stuff method, which simply stuffs all retrieved documents into a single prompt, was the fastest and most token-efficient approach but sacrificed response accuracy by not addressing query ambiguity. However, their study does not offer any insights on the energy consumption of the methods, omitting a crucial dimension for a comprehensive system evaluation. Our research puts energy consumption at the center and also studies potential trade-offs with latency and accuracy.
@@
In a study from the medical domain, 11101522 systematically evaluated the end-to-end performance of RAG systems on the RAGEval DragonBall dataset, employing retrieval metrics (recall, precision, MAP) and generation metrics (RAGAs, BERTScore) to compare cost-effective sparse and dense retrieval methods against commercial embeddings. Their key finding was that the sparse method BM25 outperformed all dense and commercial embedding models in the retrieval stage and achieved the highest scores in most generation metrics. While their research provides valuable insight into the effectiveness of open sparse retrieval methods like BM25, it is limited to a single technique and medical domain benchmark dataset. In contrast, our research extends this investigation to a real-world production application by evaluating a RAG system adopted by our industry partner. This allows us to explore the performance and challenges of these retrieval methods under practical operational conditions and constraints.
@@
Overall, there is a lack of empirical evidence about techniques that directly optimize the energy consumption in RAG systems. Despite the increasing popularity of RAG as a framework within ML-enabled applications, its energy consumption characteristics and techniques to improve them remain underexplored. Most existing studies on energy efficiency focus on general ML-enabled systems rather than RAG. While these more general techniques can still be of situational value for designing greener RAG architectures, e.g., by selecting energy-efficient algorithms (Kaack2022), more specific green RAG techniques and evidence about their effectiveness and trade-offs are needed. Our study aims to start closing this gap.
@@ 2. RQ2
To maintain a manageable scope for the trade-off analysis (RQ2), we focused on two QAs that are essential for the effectiveness and acceptable user experience of industrial RAG systems for question answering: the accuracy and latency of responses. Therefore, understanding the trade-off between energy consumption and these two QAs is crucial. Unnecessary retrieval may increase latency and computational costs, while insufficient retrieval may lead to incomplete or incorrect answers (su2025fastbetterbalancingaccuracy). Thus, investigating the trade-offs that techniques cause among these quality concerns holds significant value for practitioners and researchers alike.
@@ 3.1. Experiment Objects
T1 – Increase similarity threshold of pgvector queries: The threshold in pgvector [^2] https://github.com/pgvector/pgvector refers to the similarity score used to filter results when querying vector embeddings in PostgreSQL. This score determines which documents are returned based on their similarity to the user query. A higher threshold returns fewer, more relevant documents, thereby reducing the amount of context retrieved in a RAG system. This reduction may lead to reduced energy consumption, as less data needs to be processed. bulgakov2024optimizationretrievalaugmentedgenerationcontext showed that filtering semantically incoherent documents via thresholding significantly improves retrieval quality while minimizing memory and compute overhead. The original baseline threshold in the RAG system under study was set to 0.58. To better understand the impact of this value, we evaluated 100 test queries and found that the mean similarity score was 0.78. This provided a statistically grounded reference point for typical retrieval behavior and helped avoid arbitrary threshold selection. Since real production data is typically significantly larger and more diverse than the test sample, using the mean similarity score as a reference is more reasonable for generalization, as it captures the central tendency of similarity scores over a broader range of queries. Based on this, we chose additional thresholds at regular intervals around these two values to examine how varying similarity cutoffs affect energy consumption and the other QAs (0.58 as the baseline, 0.68, 0.78, and 0.88).
@@
T2 - Introduce lightweight reranking algorithm: Reranking typically serves as an enhancement technique to improve the accuracy of RAG systems (sharifymoghaddam2025rankllmpythonpackagereranking). Since GPU inference typically consumes more energy than CPU operations, reducing the load on the GPU can help to lower overall energy usage. In our experiment, we applied a lightweight reranking method using BM25S (bm25s) to filter and prioritize candidate documents before they are processed by the frozen LLM. This approach reduces the number of documents passed to the LLM, potentially decreasing energy consumption while maintaining performance.
@@
T3 - Reduce embedding sizes of the embedding model: Reducing the size of word embeddings can improve their efficient use in memory-constrained devices, benefiting real-world applications (raunak-etal-2019-effective). In our experiment, we replaced the default embedding model in the baseline system (e5-large-v2333https://huggingface.co/intfloat/e5-large-v2, 1024 dimensions) with smaller variants from the same provider: e5-base-v2444https://huggingface.co/intfloat/e5-base-v2 (768 dimensions) and e5-small-v2555https://huggingface.co/intfloat/e5-small-v2 (384 dimensions).
@@
T4 - Apply an efficient vector search indexing strategy: The indexing strategy of the vector search in RAG systems impacts how quickly the search is performed but also which and how many documents are selected. Previous studies indicate that indexing via Hierarchical Navigable Small World (HNSW) (HNSW_indexing) or Inverted File with Flat Quantization (IVFFlat) (IVFFlat_indexing) can enhance efficiency, particularly with binary embeddings (herranen2024sustainable). HNSW indexing constructs a graph-based structure that facilitates efficient and robust approximate nearest neighbor search by traversing hierarchical layers of nodes. IVFFlat indexing divides the data into clusters and employs inverted lists to rapidly narrow down the search space. Since both indexing methods are well-suited for fast approximate nearest neighbor searches in high-dimensional spaces and also supported by pgvector, we included them both in our experiment.
@@
T5 - Cache intermediate retrieval states via knowledge trees: If several queries show decent similarity with each other, caching the intermediate knowledge in memory for reuse can enhance system efficiency by reducing redundant computation. To allow this, jin2024ragcacheefficientknowledgecaching have proposed the RAGCache approach that uses special structures called knowledge trees. This allows new queries to skip recomputation for shared prefixes and has been shown to improve overall latency and throughput in RAG systems (lu2024turboragacceleratingretrievalaugmentedgeneration; liu2024cachegen; cheng2024large; yao2024cacheblend). We adopt a configuration in vLLM for this called enable_prefix_caching (kwon2023efficient), which enables the key-value (KV) cache of existing queries.
@@ 3.2. Experiment Materials
To ground the experiment design in real-world application requirements, we selected ChatRAG as our experiment system, a RAG system developed by SIG. ChatRAG is a query-based chatbot that enhances its responses with insights from retrieved internal documents by feeding them directly into the initial stage of the generator (zhao2024retrievalaugmentedgenerationaigeneratedcontent). The specific architecture of ChatRAG is illustrated in Fig. 1, while the detailed configuration of the baseline system and its components is provided in Table 1.
@@
To collect energy consumption data, we used Kepler [^9] https://sustainable-computing.io/, which attributes power usage to containers and Kubernetes Pods. Kepler gathers real-time power consumption metrics from node components using Intel’s Running Average Power Limit (RAPL) for CPU and DRAM power, and the NVIDIA Management Library (NVML) for GPU power. The collected and estimated container-level data is then stored using Prometheus. To host the model in a Kubernetes environment with GPU resources, we adopted vLLM (kwon2023efficientmemorymanagementlarge). vLLM is a high-throughput and memory-efficient inference and serving engine for LLMs, supporting scalable deployment on Kubernetes. With the assistance of the integration tool Helm [^10] https://helm.sh, we could efficiently deploy our models on Kubernetes.
To send and evaluate API calls to ChatRAG, we required an evaluation tool capable of measuring system performance. We adopted LLMPerf [^11] https://github.com/ray-project/llmperf.git, which supports customization for issuing calls and collecting results related to latency and accuracy. In our experiment, we extended this tool, which was originally based on the Ray distributed framework for emerging AI applications (moritz2018raydistributedframeworkemerging), to support request handling, energy data collection, and accuracy tracking. This extended version of the tool is publicly available on GitHub. [^12] https://github.com/KafkaOtto/rageval

### v2

@@ 1. Introduction
Large language models (LLMs) as the most recent wave of machine learning (ML) have shown impressive performance in natural language processing tasks and beyond (Chang et al., 2024). However, they also face several challenges, including hallucination (Ji et al., 2023; Xu et al., 2025), outdated knowledge, and opaque or untraceable reasoning processes (Gao et al., 2024). Retrieval-augmented generation (RAG) has emerged as a popular technique to address shortcomings by integrating knowledge from external databases into the generation process (Ji et al., 2023).
@@
As ML-enabled systems are increasingly adopted, more attention is also spent on their carbon footprint (Luccioni et al., 2024; Bashir et al., 2024). Due to factors such as increased technology adoption, cryptocurrency trends, and the rising demand for artificial intelligence, global data center electricity consumption is projected to range between 620 and 1,050 TWh by 2026 (International Energy Agency, 2024). The growing popularity of RAG systems is adding to this consumption, as they are increasingly applied in diverse domains such as code generation (Parvez et al., 2021; Zhou et al., 2022), question answering (Hu et al., 2022; Huang et al., 2021), and AI for science (Wang et al., 2023). Their energy consumption and carbon footprint should therefore be a subject of concern.
@@
While research on the environmental impact of ML-enabled systems has received more interest in recent years (Verdecchia et al., 2023; Gezici and Tarhan, 2022), the quality attribute (QA) of environmental sustainability has received little direct attention in RAG systems. Existing research predominantly emphasizes other QAs, such as accuracy and latency, while largely neglecting energy efficiency. For instance, Wang et al. (2024) investigated various RAG configurations but focused primarily on performance metrics like accuracy and latency. Similarly, Järvenpää et al. (2024) have proposed 30 green architectural tactics for generic ML-enabled systems, but the practical applicability of these tactics within the context of RAG systems remains largely unexplored. This underscores a clear research gap: studying how to reduce energy consumption as a core quality concern in RAG systems and systematically exploring trade-offs with other system attributes. While some techniques have been proposed to improve the resource utilization or latency of RAG systems, no systematic study has examined the energy efficiency impact of RAG techniques to provide guidance to practitioners about their usage.
@@
To close this gap, the primary objective of this research is to analyze the energy consumption of various techniques in RAG systems while also examining potential trade-offs with response latency and the accuracy of generated answers. We conducted a controlled experiment (Wohlin et al., 2012) involving five techniques, some tested under different parameter settings, resulting in a total of nine configurations that we compared to a baseline industrial RAG system. To the best of our knowledge, this is the first study to conduct an in-depth investigation into different energy techniques for RAG systems. Our findings provide valuable insights for future research on the environmental sustainability of such systems and offer guidance on navigating trade-offs between energy usage, latency, and answer accuracy.
@@ 2. Related Work
Järvenpää et al. (2024) presented a synthesis of green architectural tactics for ML-enabled systems. They compiled a catalog of 30 tactics, derived from an extensive review of Green AI literature and validated through an expert focus group. These tactics span six categories, namely data-centric strategies, algorithm design, model optimization, model training, deployment, and management, aimed at reducing energy consumption and enhancing computational efficiency throughout the ML system lifecycle. While their work provides valuable, actionable guidance for building sustainable ML-enabled systems, the proposed tactics are not specifically tailored to RAG systems, and evidence for their effectiveness in this context is missing. In this study, we focus specifically on production-level RAG systems and empirically evaluate the impact of selected techniques on energy efficiency.
@@
Using experiments to demonstrate that current LLM agents and RAG systems consume substantial amounts of energy, Wu et al. (2025) introduced the “Sustainable AI Trilemma”, highlighting the tensions between AI capability, negative environmental impact, and digital inequality. They reported that common optimization steps, e.g., query optimization or compression, drastically increase energy use with diminishing returns and that LLM-dependent methods consume orders of magnitude more energy than non-LLM alternatives. While their paper crucially highlights these issues and inefficiencies within RAG systems, it stops at diagnosis and does not propose or study specific green techniques as a solution. In contrast, our research focuses explicitly on evaluating proposed technique variations for designing green RAG systems.
@@
Current techniques targeting RAG systems primarily focus on improving other key performance metrics, such as response time or reducing computational resource usage. In several cases, this can also lead to reductions in energy consumption. For example, Arefeen et al. (2024) examined the effect of threshold k selection in chunk filtering within the iRAG framework, demonstrating that a balanced selection of k enhances query efficiency and reduces computational waste. While increasing k results in longer query processing and greater computational overhead due to processing more candidate chunks, effective chunk filtering mitigates this by minimizing unnecessary computation while preserving high recall. However, there remains a lack of broad empirical evidence on the effectiveness and trade-offs of various energy-saving techniques specifically designed for RAG systems, especially for realistic systems from an industry context.
@@
Similarly, Şakar and Emekci (2025) conducted a comprehensive evaluation of various RAG systems, identifying optimal configurations that balance critical performance metrics like response accuracy, token efficiency, runtime, and hardware utilization across diverse domains. Their findings indicate that Reciprocal RAG, which generates and ranks multiple query variations to resolve ambiguity, achieved the highest similarity score but at the cost of significantly higher token usage and longer run times. In contrast, the Stuff method, which simply stuffs all retrieved documents into a single prompt, was the fastest and most token-efficient approach but sacrificed response accuracy by not addressing query ambiguity. However, their study does not offer any insights on the energy consumption of the methods, omitting a crucial dimension for a comprehensive system evaluation. Our research puts energy consumption at the center and also studies potential trade-offs with latency and accuracy.
@@
In a study from the medical domain, Kartiyanta et al. (2025) systematically evaluated the end-to-end performance of RAG systems on the RAGEval DragonBall dataset, employing retrieval metrics (recall, precision, MAP) and generation metrics (RAGAs, BERTScore) to compare cost-effective sparse and dense retrieval methods against commercial embeddings. Their key finding was that the sparse method BM25 outperformed all dense and commercial embedding models in the retrieval stage and achieved the highest scores in most generation metrics. While their research provides valuable insight into the effectiveness of open sparse retrieval methods like BM25, it is limited to a single technique and medical domain benchmark dataset. In contrast, our research extends this investigation to a real-world production application by evaluating a RAG system adopted by our industry partner. This allows us to explore the performance and challenges of these retrieval methods under practical operational conditions and constraints.
@@
Overall, there is a lack of empirical evidence about techniques that directly optimize the energy consumption in RAG systems. Despite the increasing popularity of RAG as a framework within ML-enabled applications, its energy consumption characteristics and techniques to improve them remain underexplored. Most existing studies on energy efficiency focus on general ML-enabled systems rather than RAG. While these more general techniques can still be of situational value for designing greener RAG architectures, e.g., by selecting energy-efficient algorithms (Kaack et al., 2022), more specific green RAG techniques and evidence about their effectiveness and trade-offs are needed. Our study aims to start closing this gap.
@@ 2. RQ2
To maintain a manageable scope for the trade-off analysis (RQ2), we focused on two QAs that are essential for the effectiveness and acceptable user experience of industrial RAG systems for question answering: the accuracy and latency of responses. Therefore, understanding the trade-off between energy consumption and these two QAs is crucial. Unnecessary retrieval may increase latency and computational costs, while insufficient retrieval may lead to incomplete or incorrect answers (Su et al., 2025). Thus, investigating the trade-offs that techniques cause among these quality concerns holds significant value for practitioners and researchers alike.
@@ 3.1. Experiment Objects
T1 – Increase similarity threshold of pgvector queries: The threshold in pgvector [^2] https://github.com/pgvector/pgvector refers to the similarity score used to filter results when querying vector embeddings in PostgreSQL. This score determines which documents are returned based on their similarity to the user query. A higher threshold returns fewer, more relevant documents, thereby reducing the amount of context retrieved in a RAG system. This reduction may lead to reduced energy consumption, as less data needs to be processed. Bulgakov (2024) showed that filtering semantically incoherent documents via thresholding significantly improves retrieval quality while minimizing memory and compute overhead. The original baseline threshold in the RAG system under study was set to 0.58. To better understand the impact of this value, we evaluated 100 test queries and found that the mean similarity score was 0.78. This provided a statistically grounded reference point for typical retrieval behavior and helped avoid arbitrary threshold selection. Since real production data is typically significantly larger and more diverse than the test sample, using the mean similarity score as a reference is more reasonable for generalization, as it captures the central tendency of similarity scores over a broader range of queries. Based on this, we chose additional thresholds at regular intervals around these two values to examine how varying similarity cutoffs affect energy consumption and the other QAs (0.58 as the baseline, 0.68, 0.78, and 0.88).
@@
T2 - Introduce lightweight reranking algorithm: Reranking typically serves as an enhancement technique to improve the accuracy of RAG systems (Sharifymoghaddam et al., 2025). Since GPU inference typically consumes more energy than CPU operations, reducing the load on the GPU can help to lower overall energy usage. In our experiment, we applied a lightweight reranking method using BM25S (Lù, 2024) to filter and prioritize candidate documents before they are processed by the frozen LLM. This approach reduces the number of documents passed to the LLM, potentially decreasing energy consumption while maintaining performance.
@@
T3 - Reduce embedding sizes of the embedding model: Reducing the size of word embeddings can improve their efficient use in memory-constrained devices, benefiting real-world applications (Raunak et al., 2019). In our experiment, we replaced the default embedding model in the baseline system (e5-large-v2333https://huggingface.co/intfloat/e5-large-v2, 1024 dimensions) with smaller variants from the same provider: e5-base-v2444https://huggingface.co/intfloat/e5-base-v2 (768 dimensions) and e5-small-v2555https://huggingface.co/intfloat/e5-small-v2 (384 dimensions).
@@
T4 - Apply an efficient vector search indexing strategy: The indexing strategy of the vector search in RAG systems impacts how quickly the search is performed but also which and how many documents are selected. Previous studies indicate that indexing via Hierarchical Navigable Small World (HNSW) (Malkov and Yashunin, 2020) or Inverted File with Flat Quantization (IVFFlat) (Muja and Lowe, 2014) can enhance efficiency, particularly with binary embeddings (Herranen, 2024). HNSW indexing constructs a graph-based structure that facilitates efficient and robust approximate nearest neighbor search by traversing hierarchical layers of nodes. IVFFlat indexing divides the data into clusters and employs inverted lists to rapidly narrow down the search space. Since both indexing methods are well-suited for fast approximate nearest neighbor searches in high-dimensional spaces and also supported by pgvector, we included them both in our experiment.
@@
T5 - Cache intermediate retrieval states via knowledge trees: If several queries show decent similarity with each other, caching the intermediate knowledge in memory for reuse can enhance system efficiency by reducing redundant computation. To allow this, Jin et al. (2024) have proposed the RAGCache approach that uses special structures called knowledge trees. This allows new queries to skip recomputation for shared prefixes and has been shown to improve overall latency and throughput in RAG systems (Lu et al., 2024; Liu et al., 2024; Cheng et al., 2024; Yao et al., 2024). We adopt a configuration in vLLM for this called enable_prefix_caching (Kwon et al., 2023a), which enables the key-value (KV) cache of existing queries.
@@ 3.2. Experiment Materials
To ground the experiment design in real-world application requirements, we selected ChatRAG as our experiment system, a RAG system developed by SIG. ChatRAG is a query-based chatbot that enhances its responses with insights from retrieved internal documents by feeding them directly into the initial stage of the generator (Zhao et al., 2024). The specific architecture of ChatRAG is illustrated in Fig. 1, while the detailed configuration of the baseline system and its components is provided in Table 1.
@@
To collect energy consumption data, we used Kepler [^9] https://sustainable-computing.io/, which attributes power usage to containers and Kubernetes Pods. Kepler gathers real-time power consumption metrics from node components using Intel’s Running Average Power Limit (RAPL) for CPU and DRAM power, and the NVIDIA Management Library (NVML) for GPU power. The collected and estimated container-level data is then stored using Prometheus. To host the model in a Kubernetes environment with GPU resources, we adopted vLLM (Kwon et al., 2023b). vLLM is a high-throughput and memory-efficient inference and serving engine for LLMs, supporting scalable deployment on Kubernetes. With the assistance of the integration tool Helm [^10] https://helm.sh, we could efficiently deploy our models on Kubernetes.
To send and evaluate API calls to ChatRAG, we required an evaluation tool capable of measuring system performance. We adopted LLMPerf [^11] https://github.com/ray-project/llmperf.git, which supports customization for issuing calls and collecting results related to latency and accuracy. In our experiment, we extended this tool, which was originally based on the Ray distributed framework for emerging AI applications (Moritz et al., 2018), to support request handling, energy data collection, and accuracy tracking. This extended version of the tool is publicly available on GitHub. [^12] https://github.com/KafkaOtto/rageval